How many 5-grain tablets can be produced from 2.50 kg of powdered aspirin? (1 grain= 64.8 mg)
1
Expert's answer
2017-02-01T03:22:12-0500

Solution:
To find the answer we have to bring all masses in the task to the same units. As the base SI unit for mass is kilogram we convert masses to kilograms.
1 grain = 64.8 mg;
1 mg = 10-6 kg;

1 tablet weighs 5 grains = 5 * 64.8 mg = 324.0 mg = 324*10-6 kg = 3.24*10-4 kg

The number of tablets produced is (assuming that the number of tablets is integer)
2.50 kg / 3.24*10-4 kg = 7716 pcs.

Answer:
7716 tablets.